# Design of an Expression Recognition Solution Employing the Global Channel-Spatial Attention Mechanism

### v2

@@ Abstract
Facial expression recognition is a challenging classification task with broad application prospects in the field of human-computer interaction. This paper aims to introduce the method we will adopt in the 8th Affective and Behavioral Analysis in the Wild (ABAW) Competition, which will be held during CVPR 2025.To address the issues such as low recognition accuracy caused by subtle facial expression changes and multi-scale problems in facial expression recognition within videos, we propose a global channel-spatial attention mechanism to enhance the features initially extracted from both the speech and image modalities respectively, thereby improving the representational ability of the modalities.Secondly, in order to make full use of the complementarity between the speech and facial expression modalities, we adopt a decision fusion strategy based on the proportional criterion. We use the key frame alignment technology for speech and facial expressions to calculate the weights of speech and facial expressions, and then input these weights into the feature fusion layer for fusion, which effectively improves the recognition rate of facial expression recognition. In the facial expression recognition task of the 8th ABAW Competition, our method has achieved excellent results on the official validation set. This result fully confirms the effectiveness and competitiveness of the method we have proposed.
@@ 1 Introduction
The problem of facial expression recognition involves multiple research fields such as artificial intelligence, bioscience, and psychology, and it is a popular research branch within the field of artificial intelligence[18, 27]. Traditional facial expression recognition mainly focuses on the research of a single modality of facial images[28, 20]. However, the emotional information covered by it is limited, and different modalities of data also have different sensitivities to emotional states. Therefore, integrating multiple modalities has become a research trend. There are mainly three strategies for the integration of multiple modalities: feature-level fusion, decision-level fusion, and model fusion. Among them, the decision-level fusion method does not require strict temporal synchronization between modalities, and it also solves the problem of different feature reliabilities of different modalities. Therefore, it is a popular choice in modern emotion recognition architectures.
@@
Researchers conduct modeling and analysis of human emotions through speech - act data[42, 4] such as speech, text, and facial images, as well as physiological signal data such as electroencephalogram (EEG) and body temperature[31, 9]. Among them, the two modalities of speech and facial images account for up to 93%[6] of the conveyed emotional factors. The two are usually complementary and can provide relatively comprehensive information for the emotion recognition task. In recent years, deep - learning - related algorithms such as Convolutional Neural Network (CNN), Residual Network (ResNet), Long Short - Term Memory (LSTM)[33, 10, 48], and attention mechanisms have been widely applied to the research of speech signal feature extraction and speech emotion classification. For example, Kumaran[31] et al. proposed a deep convolutional recurrent neural network model; Chen[4] et al. proposed a Bidirectional Long and Short Term Memory (BiLSTM) network based on the attention mechanism. However, currently, dual - modality fusion models still have problems such as complex network structures, a large number of model parameters, and difficulty in training.
@@
In order to promote interdisciplinary cooperation and address key research issues spanning affective computing, machine learning, and multimodal signal processing, Kollias et al. took the lead in launching the ”Affective Behavior Analysis in the Wild” (ABAW) initiative[25]. The 8th ABAW Workshop and Competition is scheduled to be held in conjunction with the IEEE CVPR conference in 2025.
Aiming at the above emotion recognition problems, this paper conducts research based on speech and static facial images, and the main contributions are as follows:
@@
1. We propose a channel - spatial attention mechanism suitable for audio - visual emotion recognition. By combining the channel attention, channel shuffling, and spatial attention mechanisms, this mechanism aims to capture the global dependencies in the feature maps and enhance the feature representation ability of the extracted original speech and images.
@@
2. We use a proportional decision fusion strategy to obtain the final emotion classification results. A coarse - fine granularity loss function is designed to optimize the performance of the entire network.
@@
3. In the 8th ABAW Competition, our algorithm achieved excellent results on the official validation set. This result fully confirms the effectiveness and competitiveness of the proposed method.
@@ 2.1 Image-based facial expression recognition
Various visual descriptors can be extracted from facial morphology to detect expressions in video streams[24, 22]. For example, Nguyen et al.[39] extracted 68 facial landmark points and constructed 32 geometric descriptors to train a Support Vector Machine (SVM) classifier for differentiating different emotional categories. The framework proposed in reference[17, 15] is based on Facial Action Units (AUs), which continuously detects emotional states through AUs. With the development of deep learning technology, visual - based emotion recognition systems using 2D/3D Convolutional Neural Network (CNN) architectures, taking video frames or sequences as input, have higher recognition rates compared to traditional frame - aggregation - based methods. Two - dimensional convolutional neural networks like the Emotional Deep Attention Network (EmotionalDAN) [45]aim to solve the problems of emotion, valence, and landmark recognition in one step. The Spatial Transformer Network (STN)[12] can detect the main regions of interest in video frames and correct spatial variations, and a similar framework proposed in reference[29] is used to capture facial landmark points or facial visual saliency maps. Facial Expression Recognition (FER) from video streams regards a series of frames within a temporal analysis window as a single input. For example, the C3D network used in reference[46], which utilizes 3D convolutional kernels with shared weights along the time axis, has been widely applied in dynamic facial expression recognition[2, 7, 41, 53].There are also methods to design facial expression recognition models trained on single images[19]. Reference[44] proposed supervised and self - supervised learning methods to improve the classification accuracy in fine - grained and in - the - wild scenarios. The Visual Transformer with Feature Fusion (VTFF) in reference[34] can recognize emotions under extreme conditions. The Transformer - based Facial Expression Recognition (TransFER) method in reference[49] extracts rich relation - aware representations between visual descriptors. Reference[3] proposed a Graph Convolutional Network (GCN) framework to exploit the dependencies between two types of emotion recognition tasks. Abbasi et al.[1] constructed a graph - based representation for children’s facial expression recognition. Reference[37] introduced a video - based facial expression recognition method to improve the descriptiveness of embedding features using emotion - wheel information. Overall, the analysis of these advanced methods shows that it is important to distinguish between peak and non - peak video frames in video sequences. Although deep Recurrent Neural Networks/3D Convolutional Neural Networks encode the temporal dependencies of consecutive frames, their performance is not satisfactory and they are difficult to train.
@@ 2.2 Audio - integrated emotion recognition
A number of studies have applied multimodal audio-visual analysis to emotion prediction. For example, Nguyen et al.[40] integrated 3D-CNN and deep belief networks, and fused the visual and audio feature vectors using the bilinear pooling theory to recognize emotions. In reference[35], different acoustic and visual features were input into a CNN extended with an RNN, and then an average fusion was carried out at the decision-making level. Kahou et al.[14] integrated multiple deep neural networks for different data modalities to predict emotions.VAANet[54] used a specific CNN architecture combined with a late fusion strategy, but ignored the interaction between features. AVER[8], Wang et al.[50], reference[43], reference[47], etc. processed multimodal information in different ways respectively. Recently, reference[38] proposed a model that fuses audio-visual features at the model level. Hu et al.[11], MEmoBERT[52], reference[51], and multiple transformer-based frameworks[32, 5, 13] have also made their own explorations. Experiments have shown that the performance of multimodal concatenation is better.
Overall, the performance of multimodal emotion recognition is superior to that of unimodal recognition. The model-level fusion in has a good effect[38]. The importance of information in video sequences varies. Most methods adopt a two-stage shallow processing pipeline and do not make use of the complementarity between modalities. Moreover, existing methods do not consider the correlation between emotion categories, such as the polarity defined by Mikel’s emotion wheel[36]. We propose a cross - modal audio - visual fusion framework that calculates the weights of speech and images using the global channel - spatial attention mechanism and the key - frame alignment technique for speech and expressions. Except for the initial pre - training, no auxiliary data is required. Through operations such as dividing the video stream, extracting audio and images, multiple attention mechanisms are used to process features and predict emotions. At the same time, a coarse - fine granularity loss function is designed to optimize the performance of the entire network. The next section will detail this method and its modules.
@@ 3.1 Overview
The overall framework of the audio-visual emotion recognition training network in this paper is shown in Figure 1, and the entire network is implemented in an end-to-end manner. According to the processing flow of the task, it can be divided into three stages: data preprocessing, feature extraction and feature enhancement, and modality fusion and classification. Firstly, face image sequences and audio files are extracted from the original audio-visual data to achieve data preprocessing. Then, feature extraction models for face image sequences and audio sequences are designed respectively based on the residual hybrid convolutional neural network and the multi-branch convolutional neural network methods. Subsequently, the obtained sequence features are enhanced by using the global channel-spatial attention module. Finally, a proportional decision fusion strategy is used to fuse the classification results of the two single modalities to obtain an emotion probability vector, and the final emotion classification is output based on the emotion probability vector. In order to optimize the model more efficiently, a coarse-to-fine granularity loss function is designed, and the coarse granularity penalty coefficient is determined according to the emotion probability vector.
@@ 3.2 Feature Extraction
This paper employs a Multi-branch Convolutional Neural Network (MCNN) and a Residual-based Hybrid Convolutional Neural Network (RHCNN) respectively for the extraction of emotional features from audio and facial images. The MCNN structure is composed of branches for extracting features in the local, temporal, and spatial dimensions. Each different branch consists of two convolutional blocks with different receptive field sizes. In this way, receptive fields of various sizes can be utilized to fully obtain the important emotional context information in the feature maps. This not only enriches the input features but also enhances the robustness of the model against changes in emotional features. The RHCNN structure is implemented based on the concept of residual connections and depthwise separable convolutions. Firstly, the image data is input into two convolutional blocks. Each convolutional block is composed of a convolutional layer with a receptive field size of N×N, a Batch Normalization layer (BN), and a ReLU activation function layer, which are used to extract primary features and increase the nonlinearity of the features. The residual connection block is composed of a 1×1 convolution, an N×N depthwise separable convolution, and a 1×1 convolution. The 1×1 convolution block is used to change the number of channels and increase the nonlinear representation of the network, making it easier for residual connections. The receptive field size of N×N can capture more spatial information, and a channel mixing operation is used subsequently to enhance the information exchange among the channels of the feature maps.
@@ 3.3 Global Channel-Spatial Attention
We have designed a Global Channel-Spatial Attention (GCSA) module to enhance the representational ability of the input feature maps. This module combines channel attention, channel shuffling, and spatial attention mechanisms, aiming to capture the global dependencies in the feature maps.
@@ 3.4 Decision Fusion Based on the Proportion Criterion
In response to this, this paper adds a coarse - grained penalty coefficient on the basis of the cross - entropy loss while maintaining the fine - grained classification loss, and designs a coarse and fine granularity loss function, which is shown as follows:
@@ 4.1 Datasets
The 8th Workshop and Competition on Affective and Behavioral Analysis in the Wild has announced the Aff-wild2 database/[26], which is a key collection extracted from a series of studies. This resource serves as the core of the EXPR Classification Challenge[23, 30, 21, 16]. Its audio-visual dataset consists of 548 videos and approximately 2.7 million frames, which are annotated for six basic facial expressions, as well as the neutral and ”other” categories.
To enrich the depth of the dataset, we have integrated data from the AFEW and CK+ databases. The AFEW dataset is a corpus of dynamic temporal facial expression data in a near real-world environment extracted from movies. It contains 1,426 video clips in AVI format, divided into seven discrete emotions. The CK+ dataset provides data from 123 subjects, which were filmed in a laboratory environment. It includes videos of six universal expressions and the neutral expression, with the expression labels being categorical discrete values. This integration aims to enhance the comprehensiveness and practicality of our analysis. By using the Generative Adversarial Network (GAN) technology, new expression video samples are generated based on the data features of AFEW and CK+. These samples conform to the overall data distribution characteristics of Aff-wild2, thereby increasing the diversity and quantity of the data and enhancing the comprehensiveness of our dataset.
@@ 4.2 Setup
All the experiments in this paper are developed based on Python 3.10 and the PyTorch 2.2.0 deep learning framework. A GPU of the RTX 3090 model with a video memory of 24GB is used to train and test the model. The batch size is set to 16, and the model is iterated for 300 rounds. The Stochastic Gradient Descent (SGD) optimizer is selected, with a learning rate of 0.0001. The learning rate for the weight parameters of speech and facial images is set to 0.001. The Dropout in the fully - connected layer is set to 0.2, and the momentum is set to 0.9. The coarse - fine granularity loss is chosen as the loss function. The smaller the loss, the higher the prediction accuracy of the model.
@@
Base + GCSA 35.53 V
Base + GCSA + CF 39.59 V

### v3

# Design of an Expression Recognition Solution Based on the Global Channel-Spatial Attention Mechanism and Proportional Criterion Fusion

@@ Abstract
Facial expression recognition is a challenging classification task that holds broad application prospects in the field of human-computer interaction. This paper aims to introduce the method we will adopt in the 8th Affective and Behavioral Analysis in the Wild (ABAW) Competition, which will be held during the Conference on Computer Vision and Pattern Recognition (CVPR) in 2025. First of all, we apply the frequency masking technique and the method of extracting data at equal time intervals to conduct targeted processing on the original videos. Then, based on the residual hybrid convolutional neural network and the multi-branch convolutional neural network respectively, we design feature extraction models for image and audio sequences. In particular, we propose a global channel-spatial attention mechanism to enhance the features initially extracted from both the audio and image modalities respectively. Finally, we adopt a decision fusion strategy based on the proportional criterion to fuse the classification results of the two single modalities, obtain an emotion probability vector, and output the final emotional classification. We also design a coarse - fine granularity loss function to optimize the performance of the entire network, which effectively improves the accuracy of facial expression recognition. In the facial expression recognition task of the 8th ABAW Competition, our method ranked third on the official validation set. This result fully confirms the effectiveness and competitiveness of the method we have proposed.
@@ 1 Introduction
The problem of facial expression recognition involves multiple research fields such as artificial intelligence, bioscience, and psychology, and it is a popular research branch within the field of artificial intelligence. Traditional facial expression recognition mainly focuses on the research of a single modality of facial images. However, the emotional information covered by it is limited, and different modalities of data also have different sensitivities to emotional states. Therefore, integrating multiple modalities has become a research trend. There are mainly three strategies for the integration of multiple modalities: feature-level fusion, decision-level fusion, and model fusion. Among them, the decision-level fusion method does not require strict temporal synchronization between modalities, and it also solves the problem of different feature reliabilities of different modalities. Therefore, it is a popular choice in modern emotion recognition architectures.
@@
Researchers conduct modeling and analysis of human emotions through audio - act data[3] such as audio, text, and facial images, as well as physiological signal data such as electroencephalogram (EEG) and body temperature[27]. Among them, the two modalities of audio and facial images account for up to 93%[5] of the conveyed emotional factors. The two are usually complementary and can provide relatively comprehensive information for the emotion recognition task. In recent years, deep - learning - related algorithms such as Convolutional Neural Network (CNN), Residual Network (ResNet), Long Short - Term Memory (LSTM)[42], and attention mechanisms have been widely applied to the research of audio signal feature extraction and audio emotion classification. For example, Kumaran[27] et al. proposed a deep convolutional recurrent neural network model; Chen[3] et al. proposed a Bidirectional Long and Short Term Memory (BiLSTM) network based on the attention mechanism. However, currently, dual - modality fusion models still have problems such as complex network structures, a large number of model parameters, and difficulty in training.
@@
In order to promote interdisciplinary cooperation and address key research issues spanning affective computing, machine learning, and multimodal signal processing, Kollias et al. took the lead in launching the ”Affective Behavior Analysis in the Wild” (ABAW) initiative[25]. The 8th ABAW Workshop and Competition is scheduled to be held in conjunction with the IEEE CVPR conference in 2025. Aiming at the above emotion recognition problems, this paper conducts research based on audio and static facial images, and the main contributions are as follows:
@@
1.We propose a channel - spatial attention mechanism suitable for audio - visual emotion recognition. By combining the channel attention, channel shuffling, and spatial attention mechanisms, this mechanism aims to capture the global dependencies in the feature maps and enhance the feature representation ability of the extracted original audio and images.
@@
2.We use a proportional decision fusion strategy to obtain the final emotion classification results. A coarse - fine granularity loss function is designed to optimize the performance of the entire network.
@@
3.In the 8th ABAW Competition, our algorithm achieved excellent results on the official validation set. This result fully confirms the effectiveness and competitiveness of the proposed method.
@@ 2.1 Image-based facial expression recognition
Various visual descriptors can be extracted from facial morphology to detect expressions in video streams[22, 16, 20, 17]. For example, Nguyen et al.[35] extracted 68 facial landmark points and constructed 32 geometric descriptors to train a Support Vector Machine (SVM) classifier for differentiating different emotional categories. The framework proposed in reference is based on Facial Action Units (AUs), which continuously detects emotional states through AUs. With the development of deep learning technology, visual - based emotion recognition systems using 2D/3D Convolutional Neural Network (CNN) architectures, taking video frames or sequences as input, have higher recognition rates compared to traditional frame - aggregation - based methods. Two - dimensional convolutional neural networks like the Emotional Deep Attention Network (EmotionalDAN) [39]aim to solve the problems of emotion, valence, and landmark recognition in one step. The Spatial Transformer Network (STN)[8] can detect the main regions of interest in video frames and correct spatial variations, and a similar framework proposed in reference is used to capture facial landmark points or facial visual saliency maps. Facial Expression Recognition (FER) from video streams regards a series of frames within a temporal analysis window as a single input. For example, the C3D network used in reference[40], which utilizes 3D convolutional kernels with shared weights along the time axis, has been widely applied in dynamic facial expression recognition[46].There are also methods to design facial expression recognition models trained on single images. Reference[38] proposed supervised and self - supervised learning methods to improve the classification accuracy in fine - grained and in - the - wild scenarios. The Visual Transformer with Feature Fusion (VTFF) in reference[30] can recognize emotions under extreme conditions. The Transformer - based Facial Expression Recognition (TransFER) method in reference[43] extracts rich relation - aware representations between visual descriptors. Reference[2] proposed a Graph Convolutional Network (GCN) framework to exploit the dependencies between two types of emotion recognition tasks. Abbasi et al.[1] constructed a graph - based representation for children’s facial expression recognition. Reference[33] introduced a video - based facial expression recognition method to improve the descriptiveness of embedding features using emotion - wheel information. Overall, the analysis of these advanced methods shows that it is important to distinguish between peak and non - peak video frames in video sequences. Although deep Recurrent Neural Networks/3D Convolutional Neural Networks encode the temporal dependencies of consecutive frames, their performance is not satisfactory and they are difficult to train.
@@ 2.2 Audio - integrated emotion recognition
A number of studies have applied multimodal audio-visual analysis to emotion prediction. For example, Nguyen et al.[36] integrated 3D-CNN and deep belief networks, and fused the visual and audio feature vectors using the bilinear pooling theory to recognize emotions. In reference[31], different acoustic and visual features were input into a CNN extended with an RNN, and then an average fusion was carried out at the decision-making level. Kahou et al.[10] integrated multiple deep neural networks for different data modalities to predict emotions.VAANet[48] used a specific CNN architecture combined with a late fusion strategy, but ignored the interaction between features. AVER[6], Wang et al.[44], reference[37], reference[41], etc. processed multimodal information in different ways respectively. Recently, reference[34] proposed a model that fuses audio-visual features at the model level. Hu et al.[7], MEmoBERT[47], reference[4], and multiple transformer-based frameworks[28, 9] have also made their own explorations. Experiments have shown that the performance of multimodal concatenation is better.
Overall, the performance of multimodal emotion recognition is superior to that of unimodal recognition. The model-level fusion in has a good effect[34]. The importance of information in video sequences varies. Most methods adopt a two-stage shallow processing pipeline and do not make use of the complementarity between modalities. Moreover, existing methods do not consider the correlation between emotion categories, such as the polarity defined by Mikel’s emotion wheel[32]. We propose a cross-modal audio-visual fusion framework, which employs the global channel-spatial attention mechanism and the decision fusion strategy based on the proportional criterion. Except for the initial pre-training, no auxiliary data is required. Through operations such as splitting the video stream and extracting audio and images, multiple attention mechanisms are utilized to process features and predict emotions. Meanwhile, a coarse - fine granularity loss function is designed to optimize the performance of the entire network. The next section will provide a detailed introduction to this method and its modules.
@@ 3.1 Overview
The overall framework of the audio-visual emotion recognition training network in this paper is shown in Figure 1. The entire network is implemented in an end-to-end manner. According to the processing flow of the task, it can be divided into three stages: data preprocessing, feature extraction and feature enhancement, and modality fusion and classification. First is the data preprocessing stage. Face image sequences and audio files are extracted from the original audio-visual data. The frequency masking technique is applied to conduct targeted processing on the original audio data. At the same time, image data content is selected at equal time intervals to ensure the representativeness and regularity of the samples. Then, based on the residual hybrid convolutional neural network and the multi-branch convolutional neural network methods respectively, feature extraction models for face image sequences and audio sequences are designed. Subsequently, the obtained sequence features are enhanced by using the global channel-spatial attention module. Finally, a proportional decision fusion strategy is adopted to fuse the classification results of the two single modalities, obtaining an emotion probability vector, and the final emotion classification is output based on this emotion probability vector. In order to optimize the model more efficiently, a coarse - fine granularity loss function is designed, and the coarse granularity penalty coefficient is determined according to the emotion probability vector.
@@ 3.2 Feature Extraction
This paper employs a Multi-branch Convolutional Neural Network (MCNN) and a Residual-based Hybrid Convolutional Neural Network (RHCNN) respectively for the extraction of emotional features from audio and facial images.
As shown in the following formula, the audio features extracted by the MCNN are denoted as $F_{a}$, and $A$ refers to the input audio sequence. against changes in emotional features.
$F_{a}=\text{MCNN}(A)$ (1)
The MCNN structure is composed of branches for extracting features in the local, temporal, and spatial dimensions. Each different branch consists of two convolutional blocks with different receptive field sizes. In this way, receptive fields of various sizes can be utilized to fully obtain the important emotional context information in the feature maps. This not only enriches the input features but also enhances the robustness of the model
Similarly, the visual features extracted by the RHCNN are denoted as $F_{v}$, and $V$ refers to the input image sequence.
$F_{v}=\text{RHCNN}(V)$ (2)
The RHCNN structure is implemented based on the concept of residual connections and depthwise separable convolutions. Firstly, the image data is input into two convolutional blocks. Each convolutional block is composed of a convolutional layer with a receptive field size of N×N, a Batch Normalization layer (BN), and a ReLU activation function layer, which are used to extract primary features and increase the nonlinearity of the features. The residual connection block is composed of a 1×1 convolution, an N×N depthwise separable convolution, and a 1×1 convolution. The 1×1 convolution block is used to change the number of channels and increase the nonlinear representation of the network, making it easier for residual connections. The receptive field size of N×N can capture more spatial information, and a channel mixing operation is used subsequently to enhance the information exchange among the channels of the feature maps.
@@ 3.3 Global Channel-Spatial Attention
After completing the preliminary feature extraction, in order to enhance the feature representation ability, we conduct enhancement processing on the extracted audio and visual features. We have designed a Global Channel-Spatial Attention (GCSA) module to enhance the representational ability of the input feature maps. This module combines channel attention, channel shuffling, and spatial attention mechanisms, aiming to capture the global dependencies in the feature maps.
@@ 3.5 Coarse - Fine Granularity Loss
After the discrete emotion categories are coarsely grained, a qualitative analysis of the emotion categories is achieved, while the fine-grained categories are a quantitative division of the emotion categories. Theoretically, a qualitative error in an event can lead to a chain reaction of errors. Therefore, the loss caused by the misjudgment of the coarse-grained emotion categories should be amplified. In response to this, this paper adds a coarse - grained penalty coefficient on the basis of the cross - entropy loss while maintaining the fine - grained classification loss, and designs a coarse and fine granularity loss function, which is shown as follows:
@@ 4.1 Datasets
The 8th Workshop and Competition on Affective and Behavioral Analysis in the Wild has announced the Aff-wild2 database, which is a key collection extracted from a series of studies. This resource serves as the core of the EXPR Classification Challenge[24, 23, 21, 13, 11, 12, 15, 19, 14, 18, 45]. Its audio-visual dataset consists of 548 videos and approximately 2.7 million frames, which are annotated for six basic facial expressions, as well as the neutral and ”other” categories.
To enrich the depth of the dataset, we have integrated data from the AFEW and CK+ databases. The AFEW dataset is a corpus of dynamic temporal facial expression data in a near real-world environment extracted from movies[26]. It contains 1,426 video clips in AVI format, divided into seven discrete emotions. The CK+ dataset provides data from 123 subjects, which were filmed in a laboratory environment[29]. It includes videos of six universal expressions and the neutral expression, with the expression labels being categorical discrete values. This integration aims to enhance the comprehensiveness and practicality of our analysis. By using the Generative Adversarial Network (GAN) technology, new expression video samples are generated based on the data features of AFEW and CK+. These samples conform to the overall data distribution characteristics of Aff-wild2, thereby increasing the diversity and quantity of the data and enhancing the comprehensiveness of our dataset.
@@ 4.2 Setup
All the experiments in this paper are developed based on Python 3.10 and the PyTorch 2.2.0 deep learning framework. A GPU of the RTX 3090 model with a video memory of 24GB is used to train and test the model. The batch size is set to 16, and the model is iterated for 300 rounds. The Stochastic Gradient Descent (SGD) optimizer is selected, with a learning rate of 0.0001. The learning rate for the weight parameters of audio and facial images is set to 0.001. The Dropout in the fully - connected layer is set to 0.2, and the momentum is set to 0.9. The coarse - fine granularity loss is chosen as the loss function. The smaller the loss, the higher the prediction accuracy of the model.
@@
Base + GCSA 33.53 V
Base + GCSA + CF 35.59 V